• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name cisbani

    Log entry time 08:03:31 on May 3,2011

    Entry number 357940

    keyword=Owl Shift Summary

    Owl Shift Summary May/03/2011

    00:00 Run 4089 ongoing in 3He 20 cm, 115 uA, kin 3.1 left, 6.5 right

    00:05 ROC10 not responding, coda restarted, run 4089 is junk

    00:30 Move to carbon target

    00:38 Target on 12C

    01:05 Ask MCC to increase the ramp up of the current: was 2 uA/s move to 5 uA/s (as in Hall C)

    02:15 Move to 40Ca

    02:20 Ask MCC 40 uA and raster size 6x4 (MCC unit), was 4.5x3

    03:30 ROC10 disconnected, coda rebooted, run affected 4096

    05:45 Move to Dummy

    05:50 Ask MCC 40 uA, raster 4.5x3

    07:05 Move target to 48Ca

    07:10 Ask MCC 40 uA, raster 6x4

    Statistics
    Run Left Cumulated
    4090 334k 3600k # Last run with 3He
    -- change target to 12 C
    4091 868k 868k # First Run with 12C 5mm / 5 mA
    4092 869k 1737k
    4093 873k 2610k
    4094 873k 3482k
    --- change target to 40 Ca
    4095 636k 636k
    4096 407k 1043k # coda restarted
    4097 637k 1680k
    4098 626k 2306k # coda frozen after run end, restarted
    4099 615k 2921k # coda frozen after run end,
    --- change target to Dummy 20 cm
    4100 0.5 h
    4101 0.5 h
    --- change target to 48 Ca
    4102 682k 682k
    4103 see next shift